# Moment maps and stability of holomorphic submersions

### v1

@@ 2.3. Deformation theory of Kähler metrics with constant scalar curvature
The above proof of Theorem 2.10 gives a different perturbation of the Kähler structure than the one given by Székelyhidi [43] and Brönnle [2], to ensure that the scalar curvature belongs to the Lie algebra of $K$. In fact, they fix the Kähler form $\omega$ and perturb the Kuranishi map; we instead perturb the Kähler form, thus the Kuranishi map remains holomorphic.
Until this point, we have considered deformations of the complex structure of a fixed symplectic manifold as $(0,1)$-forms with values in the $(1,0)$-tangent bundle. The point of view which we now adopt following Dervan–Hallam [10], consists of considering deformations of the manifold $M$ as a smooth proper morphism $\mathcal{U}\to V$, called the universal family, such that $\mathcal{U}$ is diffeomorphic to $M\times V$ and the morphism is equivariant with respect to the action of $K$.
Let $\omega_{\mathcal{U}}$ be the pull-back to $\mathcal{U}$ of the Kähler form $\omega$ of $M$; $\mathcal{U}$ admits a complex structure $\mathbb{J}$ such that $\mathbb{J}|_{x}=\Phi(x)$. The two-form $\omega_{\mathcal{U}}$ is relatively Kähler with respect to $\mathbb{J}$. There exists a moment map $\tau:\mathcal{U}\to\mathfrak{k}$ with respect to $\omega$. Even if $\omega$ is only relatively Kähler, one can still define a moment map by requiring that $\tau$ is equivariant and that for each $\xi\in\mathrm{Lie}(K)$
$\mathrm{d}\langle\tau,\xi\rangle=\omega(\cdot,\sigma_{\xi}),$
where $\sigma_{\xi}$ is the infinitesimal vector field associated to $\xi$ [10, Definition 3.3].
The relatively Kähler form $\omega_{\mathcal{U}}$ induces the two-form on $V$
$\Omega_{WP}=\frac{\widehat{S}}{n+1}\int_{\mathcal{U}/V}\omega_{\mathcal{U}}^{n+1}-\int_{\mathcal{U}/V}\rho_{\mathcal{U}}\wedge\omega_{\mathcal{U}}^{n},$
where $\rho_{\mathcal{U}}$ is the curvature of the Hermitian connection induced by $\omega_{\mathcal{U}}$ on the top wedge power of the vertical tangent bundle. It follows form Fujiki–Schumacher [20, §7] that $\Omega_{WP}$ is a Kähler metric on $V$, called the Weil–Petersson metric, and it is in fact the pull-back via $\Phi$ of the Kähler metric on $\mathscr{J}$ induced by the Hermitian inner product
$\langle\alpha,\beta\rangle_{J}:=\int_{M}\langle\alpha,\beta\rangle_{g_{J}}\frac{\omega^{n}}{n!},\quad\alpha,\beta\in T_{J}\mathscr{J}.$
The following interpretation of the scalar curvature as a moment map is due to Dervan–Hallam [10, Theorem 4.6].
@@
A moment map for the $K$-action on $V$ with respect to $\Omega_{WP}$ is given by the projection
$\langle\mu(x),\xi\rangle=\int_{\mathcal{U}_{x}}\langle\tau,\xi\rangle|_{x}\left(\mathrm{Scal}(\omega,\Phi(x))-\widehat{S}\right)\omega^{n}.$ (2.7)
In particular, Theorem 2.12 applies to any relatively Kähler metric on $\mathcal{U}\to V$. There is a difference in sign with [10]: this is due to the fact that we take the opposite sign for the infinitesimal action (2.3). Moreover, the Weil–Petersson metric and Theorem 2.12 in principle only work on the complex subspace of $V$ that parametrises integrable complex structures, but we can extend $\Omega_{WP}$ to $V$, so that $\mu$ can be extended as well in a unique way.
As before, let us denote by
@@
    $\mathfrak{k}^{\mathcal{U}}$ the Lie algebra of $K$ realised as a space of fibrewise holomorphy potentials on $\mathcal{U}$ with respect to $\omega_{\mathcal{U}}$;
@@
    $\mathfrak{k}^{\mathcal{U}}_{\varphi}$ the Lie algebra of $K$ realised as a space of fibrewise holomorphy potentials on $\mathcal{U}$ with respect to $\omega_{\mathcal{U}}+i\partial\bar{\partial}\varphi$, for a Kähler potential $\varphi$,
and analogously for $\mathfrak{k}^{V}$. Moreover, for each $\xi\in\mathfrak{k}$, we denote the induced vector fields $\sigma_{\xi}^{\mathcal{U}}$ on $\mathcal{U}$, $\sigma_{\xi}^{M}$ on $M$ and $\sigma_{\xi}^{V}$ on $V$.
Combinig Theorem 2.10 with Lemma 2.12, we obtain the following result, whose proof is identical to the one of [40, Lemma 4.15], but we explicitly adapt it here to the context of cscK metrics for sake of completeness.
@@
A point $x\in V$ is a zero of the moment map $\mu$ if and only if the scalar curvature $\mathrm{Scal}(\omega+i\partial\bar{\partial}\varphi_{x},\Phi(x))$ is constant.
@@
Let $\mathrm{pr}_{M}:\mathcal{U}\to M$ be the projection. From the second item of Theorem 2.10, for every $x\in V$ we can change $\omega_{\mathcal{U}}$ by a Kähler potential $\varphi_{x}$ with respect to the complex structure $\Phi(x)$, in such a way that
@@
Since $\varphi_{x}$ varies smoothly with $x$, it defines a relative Kähler potential $\varphi$ on $\mathcal{U}$, so the condition (2.8) can be rephrased as
$\mathrm{Scal}((\omega_{\mathcal{U}}+i\partial\bar{\partial}\varphi_{,}\mathbb{J})|_{\mathcal{U}_{x}})-\widehat{S}\in\mathrm{pr}_{M}^{*}(\mathfrak{k}^{M}_{\varphi_{x}}).$
Applying Theorem 2.12, we obtain a moment map $\mu$ on $V$ induced by the relatively Kähler metric $(\omega_{\mathcal{U}}+i\partial\bar{\partial}\varphi,\mathbb{J})$.
Next, for each $x\in V$ we define two immersions of $\mathfrak{k}$ in $C^{\infty}(M)$:
$\iota_{1,x}(\xi)=$ $(h^{\mathcal{U}}_{\xi}+\sigma_{\xi}^{\mathcal{U}}(\varphi_{x}))|_{U_{x}}$
$\iota_{2,x}(\xi)=$ $h_{\xi}+\sigma_{\xi}^{M}(\varphi_{x})$
The two immersions are injective linear maps, hence their images are isomorphic to $\mathfrak{k}$ and are given by $\iota_{1,x}(\mathfrak{k})=\mathfrak{k}_{\varphi}^{\mathcal{U}}|_{\mathcal{U}_{x}}$, $\iota_{2,x}(\mathfrak{k})=\mathfrak{k}_{\varphi_{x}}^{M}$. We can rephrase the condition (2.8) as
@@
On the other hand, Theorem 2.12 implies that a point $x\in V$ is a zero of the moment map $\mu$ if and only if the projection of the scalar curvature of $(\omega+i\partial\bar{\partial}\varphi_{x},\Phi(x))$ is $L^{2}(g_{x})$-orthogonal to $\mathfrak{k}_{\varphi}^{\mathcal{U}}|_{\mathcal{U}_{x}}$ in $C^{\infty}(M)$. Denoting by $\Pi_{1,x}$ said orthogonal projection, Theorem 2.12 implies that $\mu(x)=0$ if and only if
@@
Let $\omega_{\mathcal{U}}$ be the relatively Kähler metric on $\mathcal{U}$ after the change in the holomorphy potential $\varphi$ coming from Theorem 2.10. We next compute an expansion of the induced moment map $\mu$ about the origin of $V$. By definition,
@@
The map $\nu$ can be characterised as a moment map and is related to the scalar curvature (2.7) as follows [26, §3]:
@@ 3. Optimal symplectic connections
A relative version of Ehresmann’s theorem [39, Proposition 4.5] implies that $X$ and $Y$ are diffeomorphic. Let $M$ denote the underlying smooth manifold. Since the Chern classes are integral classes we have that $c_{1}(H_{X})$ coincides with $c_{1}(H_{Y})$ as cohomology classes on $M$. Since $Y$ is a small deformation of $X$, the cohomology class $c_{1}(H_{X})$ is of type $(1,1)$ on $Y$, so $\omega$ is a $(1,1)$-form with respect to the complex structure of $Y$ [25, §6.1]. By Moser’s theorem [3, Theorem 7.2], we can modify the complex structure of $Y$ by a small diffeomorphism so that $\omega$ restricted to the fibres of $\pi_{Y}$ is compatible with the restriction of the complex structure. Thus we can assume that $\omega$ is relatively Kähler on $Y$.
@@ 3.1. Splitting of the function space
Consider the real vector bundle $E\to B$ [12, §3.1], whose fibre over $b\in B$ is the real finite-dimensional vector space $\mathrm{ker}_{0}(\mathcal{D}_{b}^{*}\mathcal{D}_{b})$ of holomorphy potentials on the fibre $X_{b}$ with mean-value zero with respect to $\omega_{b}$. It is well defined as a vector bundle since we assume that the complex dimension of the space $H^{0}(X_{b},T^{1,0}X_{b})$ of holomorphic vector fields on $X_{b}$ is independent of $b$ [23, §2.3]. The space of smooth global sections of $E$, denoted by $C^{\infty}(E)$, is given by the kernel over the fibrewise mean-value-zero functions of the vertical Lichnerowicz operator $\mathcal{D}_{\mathcal{V}}^{*}\mathcal{D}_{\mathcal{V}}$. In [23, Lemma 2.7], Hallam used the Cartan decomposition for the space $\mathfrak{h}(X_{b})$ of holomorphic vector fields of the fibre to show that $E_{b}$ can be also viewed as the vector space of all Kähler potentials $\varphi_{b}$ on $X_{b}$ of mean-value zero for which $\omega_{b}+i\partial\bar{\partial}\varphi_{b}$ is still cscK.
@@ 3.2. Optimal symplectic connections
which is the relative version of the map (2.3). Let $\widetilde{H}^{1}_{\mathcal{V}}$ be the kernel of the elliptic [39, §4.2] operator
@@
The group $K_{\pi}$ acts on $\widetilde{H}^{1}_{\mathcal{V}}$ by pull-back. The following theorem [39, §4.2] is a fibrewise version of Theorem 2.10.
@@
[39, §3.3] A relatively Kähler metric $\omega$ on $Y\to B$ is called an optimal symplectic connection if
@@
Equation (3.4) is a second-order elliptic equation on the vector bundle $E\to B$ [39, §5.3]. It arises as the subleading order term in the expansion of the scalar curvature, as shown by the following proposition [39, Proposition 5.4].
@@
The linearisation of the equation at a solution is given by the operator [39, §5.3]
@@
The adjoint is computed with respect to $\omega_{F}+\omega_{B}$. Here $\nabla_{\mathcal{V}}^{1,0}\varphi_{E}$ is a section of the holomorphic tangent bundle; the vertical part of ¯$\bar{\partial}\nabla_{\mathcal{V}}^{1,0}\varphi_{E}$ vanishes since $\varphi_{E}\in C^{\infty}(E)$ and the horizontal part is denoted by the expression (3.5). The operator (3.5) can be described as follows [12, §4.3]: let $\mathcal{D}_{k}^{*}\mathcal{D}_{k}$ be the Lichnerowicz operator with respect to the Kähler metric $\omega_{k}$. It admits a power series expansion in negative powers of $k$:
@@
where $\varphi,\psi\in C^{\infty}(E)$. A function $\psi\in C^{\infty}(E)$ is in the kernel of $A$ if and only if $\psi$ is a fibrewise holomorphy potential with respect to all $J_{s}$, i.e. $\psi\in C^{\infty}(E,J_{s})$. Thus the kernel of $\widehat{\mathcal{L}}$ consists of those functions $\psi\in C^{\infty}(E,J_{0})$ such that ¯s($\bar{\partial}_{s}(\nabla_{s,\mathcal{V}}^{1,0}\psi)=0$ for all $s$ [39, Proposition 5.8].

### v2

@@ 2.3. Deformation theory of Kähler metrics with constant scalar curvature
The above proof of Theorem 2.10 gives a different perturbation of the Kähler structure than the one given by Székelyhidi [46] and Brönnle [3], to ensure that the scalar curvature belongs to the Lie algebra of $K$. In particular, they fix the Kähler form $\omega$ and perturb the Kuranishi map, which leads them to differentiate the function $\mathrm{Scal}(\omega,\Phi(x))$ in the $x$-variable. We instead perturb the Kähler form, thus the Kuranishi map remains holomorphic, and we differentiate the scalar curvature in the direction of the Kähler potential.
Until this point, we have considered deformations of the complex structure of a fixed symplectic manifold as $(0,1)$-forms with values in the $(1,0)$-tangent bundle. The point of view which we now adopt following Dervan–Hallam [11], consists of considering deformations of the manifold $M$ as a smooth proper morphism $\mathcal{U}\to V$, called the universal family, such that $\mathcal{U}$ is diffeomorphic to $M\times V$ and the morphism is equivariant with respect to the action of $K$. The fibre over $x\in V$ is given by the almost complex manifold $\mathcal{U}_{x}=(M,\omega,\Phi(x))$.
Let $\omega_{\mathcal{U}}$ be the pull-back to $\mathcal{U}$ of the 2-form $\omega$ of $M$; $\omega_{\mathcal{U}}$ is symplectic when restricted to the fibres and Kähler when restricted to the integrable fibres. The universal family $\mathcal{U}$ admits an almost complex structure $\mathbb{J}$ such that $\mathbb{J}|_{x}=\Phi(x)$. The two-form $\omega_{\mathcal{U}}$ is closed and relatively symplectic, and there exists a moment map $\tau:\mathcal{U}\to\mathfrak{k}$ with respect to $\omega_{\mathcal{U}}$: even if $\omega_{\mathcal{U}}$ is only relatively symplectic, one can still define the moment map by requiring that $\tau$ is equivariant and that for each $\xi\in\mathfrak{k}$
$\mathrm{d}\langle\tau,\xi\rangle=\omega_{\mathcal{U}}(\cdot,\sigma_{\xi}),$
where $\sigma_{\xi}$ is the infinitesimal vector field associated to $\xi$ [11, Definition 3.3].
The relatively symplectic form $\omega_{\mathcal{U}}$ induces the two-form on $V$
$\Omega_{WP}=\frac{\widehat{S}}{n+1}\int_{\mathcal{U}/V}\omega_{\mathcal{U}}^{n+1}-\int_{\mathcal{U}/V}\rho_{\mathcal{U}}\wedge\omega_{\mathcal{U}}^{n},$ (2.7)
where $\rho_{\mathcal{U}}$ is the curvature of the Hermitian connection induced by $\omega_{\mathcal{U}}$ on the top wedge power of the vertical tangent bundle. It follows form Fujiki–Schumacher [22, §7] that $\Omega_{WP}$ is a Kähler metric on $V$, called the Weil–Petersson metric. In fact, the Weil-Petersson metric is the pull-back via $\Phi$ of the Kähler metric on $\mathscr{J}$ induced by the Hermitian inner product
$\langle\alpha,\beta\rangle_{J}:=\int_{M}\langle\alpha,\beta\rangle_{g_{J}}\frac{\omega^{n}}{n!},\quad\alpha,\beta\in T_{J}\mathscr{J},$
which is the Donaldson-Fujiki Kähler metric on $\mathscr{J}$ [17, 21] (see also [47, §6.1]). The following interpretation of the scalar curvature as a moment map is due to Dervan–Hallam [11, Theorem 4.6].
@@
There exists a moment map $\mu$ for the $K$-action on $V$ with respect to $\Omega_{WP}$, which, on the smooth locus of $V^{\mathrm{int}}$, is given as the projection
$\langle\mu(x),\xi\rangle=\int_{\mathcal{U}_{x}}\langle\tau,\xi\rangle|_{x}\left(\mathrm{Scal}(\omega,\Phi(x))-\widehat{S}\right)\omega^{n},$ (2.8)
for $\xi\in\mathfrak{k}$.
There is a difference in sign with [11]: this is due to the fact that we take the opposite sign for the infinitesimal action (2.3). Moreover, since the smooth locus of $V^{\mathrm{int}}$ is dense in $V^{\mathrm{int}}$, and the (Chern) scalar curvature is defined for all $x\in V$, the expression (2.8) actually holds on all $V^{\mathrm{int}}$.
We now change $(\omega,J_{x})$ to $(\omega_{\varphi_{x}},J_{x})$ using Theorem 2.10 and, since $\varphi_{x}$ vary smoothly with $x$, we also change $\omega_{\mathcal{U}}$ to $\omega_{\mathcal{U},\varphi}$ globally on $\mathcal{U}$. Consider the function $\mathcal{M}_{\omega_{\mathcal{U}},\varphi}:V\to\mathbb{R}$ defined as
$\mathcal{M}_{\omega_{\mathcal{U}},\varphi}(x)=\int_{0}^{1}\int_{\mathcal{U}_{x}}\varphi_{x}\left(\mathrm{Scal}(\omega_{t\varphi_{x}},\Phi(x))-\widehat{S}\right)\omega_{t\varphi_{x}}^{n}\mathrm{d}t,$
where $(\omega_{t\varphi_{x}},J_{x})=(\omega+\frac{1}{2}(\mathrm{d}\mathrm{d}^{c}(t\varphi_{x}))^{1,1},J_{x})$. This is a smooth function on $V$, and, when $x\in V^{\mathrm{int}}$, $\mathcal{M}_{\omega_{\mathcal{U}},\varphi}(x)$ coincides with (minus) the Mabuchi functional of $\varphi_{x}$ on the corresponding fibre $\mathcal{U}_{x}$. On the integrable locus, where both $\omega_{\mathcal{U}}$ and $\omega_{\mathcal{U},\varphi}$ are relatively Kähler, the expression (2.7) defines the Weil-Petersson metrics $\Omega_{WP}$ and $\Omega_{WP,\varphi}$. Then we have the following relation [43, §7], [13, Proposition 4.2].
Proposition 2.13.
On the smooth locus of $V^{\mathrm{int}}$,
$\Omega_{WP}-\Omega_{WP,\varphi}=-i\partial\bar{\partial}\mathcal{M}_{\omega_{\mathcal{U}},\varphi}.$
Since the function $\mathcal{M}_{\omega_{\mathcal{U}},\varphi}$ is a smooth function defined on all of $V$, we can use Proposition 2.13 to extend the Weil-Petersson metric obtained using $\omega_{\mathcal{U},\varphi}$. In fact, although the expression (2.7) is only valid on the integrable locus, we define
$\Omega_{WP,\varphi}=\Omega_{WP}+i\partial\bar{\partial}\mathcal{M}_{\omega_{\mathcal{U}},\varphi},$
which is a Kähler metric defined on all of $V$ (possibly after shrinking $V$). Moreover, in the proof of Theorem 2.10, we can restrict ourselves to work only with $K$-invariant potentials, so that the $K$-equivariant implicit function theorem guarantees that $\varphi$ is $K$-invariant; hence $\mathcal{M}_{\omega_{\mathcal{U},\varphi}}$ is $K$-invariant. Thus the moment map defined by Theorem 2.12 with respect to the Weil-Petersson metric $\Omega_{WP,\varphi}$, which has the expression (2.8) on $V^{\mathrm{int}}$, extends continuously to a moment map $\mu_{\varphi}$ on all $V$ as
$\langle\mu_{\varphi},\xi\rangle=\langle\mu,\xi\rangle+\mathrm{d}^{c}\mathcal{M}_{\omega_{\mathcal{U},\varphi}}(\sigma_{\xi}).$
The extension is unique up to elements in the centre of $\mathfrak{k}$. This is the moment map we will work with for the rest of this section and in Appendix A; in the following, we denote the moment map $\mu_{\varphi}$ again by $\mu$, making the change in the potential $\varphi$ implicit.
Let us denote by
@@
    $\mathfrak{k}^{M}$ the realisation of $\mathfrak{k}$ as the space of holomorphy potentials with respect to $(\omega,J)$;
@@
Moreover, for each $\xi\in\mathfrak{k}$, we denote the induced vector fields $\sigma_{\xi}^{\mathcal{U}}$ on $\mathcal{U}$, $\sigma_{\xi}^{M}$ on $M$ and $\sigma_{\xi}^{V}$ on $V$. Combining Theorem 2.10 with Lemma 2.12, we obtain the following result, whose proof is identical to the one of [42, Lemma 4.15], but we explicitly adapt it here to the context of cscK metrics for sake of completeness.
@@
A point $x\in V^{\mathrm{int}}$ is a zero of the moment map $\mu$ if and only if the scalar curvature $\mathrm{Scal}(\omega+i\partial\bar{\partial}\varphi_{x},\Phi(x))$ is constant.
@@
Let $\mathrm{pr}_{M}:\mathcal{U}\to M$ be the projection. From the second item of Theorem 2.10, for every $x\in V^{\mathrm{int}}$ we can change $\omega$ by a Kähler potential $\varphi_{x}$ with respect to the complex structure $\Phi(x)$, in such a way that
@@
Since $\varphi_{x}$ varies smoothly with $x$, it defines a function $\varphi$ on $\mathcal{U}$, so the condition (2.9) can be rephrased as
$\mathrm{Scal}((\omega_{\mathcal{U},\varphi},\mathbb{J})|_{\mathcal{U}_{x}})-\widehat{S}\in\mathrm{pr}_{M}^{*}(\mathfrak{k}^{M}_{\varphi_{x}}).$
Next, we define two immersions of $\mathfrak{k}$ in $C^{\infty}(M)$:
$\iota_{1,x}(\xi)=$ $(h^{\mathcal{U}}_{\xi}+\sigma_{\xi}^{\mathcal{U}}(\varphi))|_{U_{x}}$
$\iota_{2,x}(\xi)=$ $h_{\xi}+\sigma_{\xi}^{M}(\varphi_{x}).$
The two immersions are injective linear maps, hence their images are isomorphic to $\mathfrak{k}$. We can rephrase the condition (2.9) as
@@
On the other hand, Theorem 2.12 implies that a point $x\in V^{\mathrm{int}}$ is a zero of the moment map $\mu$ if and only if the projection of the scalar curvature of $(\omega+i\partial\bar{\partial}\varphi_{x},\Phi(x))$ is $L^{2}(g_{x})$-orthogonal to $\iota_{1,x}(\mathfrak{k})$ in $C^{\infty}(M)$. Denoting by $\Pi_{1,x}$ said orthogonal projection, Theorem 2.12 implies that $\mu(x)=0$ if and only if
@@
Let $\omega_{\mathcal{U}}$ be the relatively Hermitian metric on $\mathcal{U}$ after the change given by the potential $\varphi$ coming from Theorem 2.10. We next compute an expansion of the induced moment map $\mu$ about the origin of $V$. By definition,
@@
The map $\nu$ can be characterised as a moment map and is related to the scalar curvature (2.8) as follows [28, §3]:
@@ 3. Optimal symplectic connections
A relative version of Ehresmann’s theorem [41, Proposition 4.5] implies that $X$ and $Y$ are diffeomorphic. Let $M$ denote the underlying smooth manifold. Since the Chern classes are integral classes we have that $c_{1}(H_{X})$ coincides with $c_{1}(H_{Y})$ as cohomology classes on $M$. Since $Y$ is a small deformation of $X$, the cohomology class $c_{1}(H_{X})$ is of type $(1,1)$ on $Y$, so $\omega$ is a $(1,1)$-form with respect to the complex structure of $Y$ [27, §6.1]. By Moser’s theorem [4, Theorem 7.2], we can modify the complex structure of $Y$ by a small diffeomorphism so that $\omega$ restricted to the fibres of $\pi_{Y}$ is compatible with the restriction of the complex structure. Thus we can assume that $\omega$ is relatively Kähler on $Y$.
@@ 3.1. Splitting of the function space
Consider the real vector bundle $E\to B$ [14, §3.1], whose fibre over $b\in B$ is the real finite-dimensional vector space $\mathrm{ker}_{0}(\mathcal{D}_{b}^{*}\mathcal{D}_{b})$ of holomorphy potentials on the fibre $X_{b}$ with mean-value zero with respect to $\omega_{b}$. It is well defined as a vector bundle since we assume that the complex dimension of the space $H^{0}(X_{b},T^{1,0}X_{b})$ of holomorphic vector fields on $X_{b}$ is independent of $b$ [25, §2.3]. The space of smooth global sections of $E$, denoted by $C^{\infty}(E)$, is given by the kernel over the fibrewise mean-value-zero functions of the vertical Lichnerowicz operator $\mathcal{D}_{\mathcal{V}}^{*}\mathcal{D}_{\mathcal{V}}$. In [25, Lemma 2.7], Hallam used the Cartan decomposition for the space $\mathfrak{h}(X_{b})$ of holomorphic vector fields of the fibre to show that $E_{b}$ can be also viewed as the vector space of all Kähler potentials $\varphi_{b}$ on $X_{b}$ of mean-value zero for which $\omega_{b}+i\partial\bar{\partial}\varphi_{b}$ is still cscK.
@@ 3.2. Optimal symplectic connections
which is the relative version of the map (2.3). Let $\widetilde{H}^{1}_{\mathcal{V}}$ be the kernel of the elliptic [41, §4.2] operator
@@
The group $K_{\pi}$ acts on $\widetilde{H}^{1}_{\mathcal{V}}$ by pull-back. The following theorem [41, §4.2] is a fibrewise version of Theorem 2.10.
@@
[41, §3.3] A relatively Kähler metric $\omega$ on $Y\to B$ is called an optimal symplectic connection if
@@
Equation (3.4) is a second-order elliptic equation on the vector bundle $E\to B$ [41, §5.3]. It arises as the subleading order term in the expansion of the scalar curvature, as shown by the following proposition [41, Proposition 5.4].
@@
The linearisation of the equation at a solution is given by the operator [41, §5.3]
@@
The adjoint is computed with respect to $\omega_{F}+\omega_{B}$. Here $\nabla_{\mathcal{V}}^{1,0}\varphi_{E}$ is a section of the holomorphic tangent bundle; the vertical part of ¯$\bar{\partial}\nabla_{\mathcal{V}}^{1,0}\varphi_{E}$ vanishes since $\varphi_{E}\in C^{\infty}(E)$ and the horizontal part is denoted by the expression (3.5). The operator (3.5) can be described as follows [14, §4.3]: let $\mathcal{D}_{k}^{*}\mathcal{D}_{k}$ be the Lichnerowicz operator with respect to the Kähler metric $\omega_{k}$. It admits a power series expansion in negative powers of $k$:
@@
where $\varphi,\psi\in C^{\infty}(E)$. A function $\psi\in C^{\infty}(E)$ is in the kernel of $A$ if and only if $\psi$ is a fibrewise holomorphy potential with respect to all $J_{s}$, i.e. $\psi\in C^{\infty}(E,J_{s})$. Thus the kernel of $\widehat{\mathcal{L}}$ consists of those functions $\psi\in C^{\infty}(E,J_{0})$ such that ¯s($\bar{\partial}_{s}(\nabla_{s,\mathcal{V}}^{1,0}\psi)=0$ for all $s$ [41, Proposition 5.8].